a patient in a hospital is often administered an intravenous (IV) drip contains an aqueous solution, this aqueous solution contains 0.85% (mass by volume) of sodium chloride or 5% (mass by volume) of glucose. calculate the molarity of both these solution.


1
Expert's answer
2022-04-17T11:04:25-0400

0.85 g NaCl in 100 mL (100 g approximately)

0.85 / 58.44 = 0.0145 (mol)

V = 100 mL = 0.1 L

0.0145 / 0.1 = 0.145 M

 

5 g glucose in 100 g (100 mL)

5 / 180.156 = 0.0278 (mol)

V = 100 mL = 0.1 L

0.0278 / 0.1 = 0.278 M